### v5

# REINFORCE++: An Efficient RLHF Algorithm with Robustness to Both Prompt and Reward Models

Abstract
Reinforcement Learning from Human Feedback (RLHF) is crucial in aligning large language models (LLMs) with human values and preferences. While state-of-the-art applications like ChatGPT/GPT-4 commonly employ Proximal Policy Optimization (PPO), including a critic network introduces significant computational overhead. REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network. However, these approaches face challenges in accurate advantage estimation. Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on more straightforward prompts and vulnerability to reward hacking. To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the normalized reward of a batch as the baseline. Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation. Furthermore, it achieves superior generalization in RLHF and long chain-of-thought (CoT) settings compared to REINFORCE-based methods. The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
1 Introduction
Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning large language models (LLMs) with human values and preferences  (Vemprala et al., 2023; Achiam et al., 2023; Ouyang et al., 2022a; Shen & Zhang, 2024). Despite the emergence of non-RL alternatives like DPO  (Rafailov et al., 2023), state-of-the-art applications such as ChatGPT/GPT-4  (Vemprala et al., 2023; OpenAI, 2023), Claude  (Anthropic, 2023), and Gemini  (Team et al., 2023) continue to rely on RL algorithms, particularly PPO, for policy optimization. However, PPO  (Schulman et al., 2017) requires a critic network, introducing substantial computational overhead and memory demands that limit large model alignment in small-scale clusters. To address this, researchers have proposed various REINFORCE-based methods that eliminate the critic network, including ReMax  (Li et al., 2023), REINFORCE Leave One-Out (RLOO)  (Wu et al., 2024), and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024). Furthermore, Deepseek-R1 demonstrates the effectiveness of the REINFORCE-based method in the long-form CoT setting  (Guo et al., 2025), which achieved state-of-the-art performance on challenging datasets using GRPO with rule-based rewards.
Without the critic network, REINFORCE-based methods often struggle to estimate individual tokens’ advantages accurately. Various REINFORCE-baseline approaches have been proposed to address this limitation, but each has significant drawbacks. ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation. RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO utilizes a normalized reward across all responses. These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking. Furthermore, these methods calculate reward baselines separately for each prompt, leading to overfitting and instability on specific training prompts during optimization. Therefore, these methods require carefully curating prompt sets particular to each task.
To address these challenges, we propose REINFORCE++, a novel REINFORCE-based method that eliminates the critic model from PPO and uses the mean reward of a global batch as the baseline. This approach prevents overfitting to specific training prompts and demonstrates robustness across both Bradley-Terry and rule-based reward models. Notably, REINFORCE++ eliminates the need for prompt set truncation and achieves strong generalization performance in both RLHF and long CoT RL settings.
The remainder of this paper is organized as follows. In Section 2, we analyze the differences between various REINFORCE-based RLHF methods, demonstrating the ineffectiveness of prompt-specific reward baselines and the overfitting problems in RLOO and GRPO methods. Section 3 presents REINFORCE++, detailing its implementation and advantages. In Section 4, we evaluate REINFORCE++ through comprehensive experiments. We first compare REINFORCE++ with other RLHF methods using both Bradley-Terry and Rule-Based Reward Models, showing its superior or comparable performance. Then, under the long-form Chain-of-Thought (CoT) setting, we focus on comparing REINFORCE++ with GRPO, the current state-of-the-art REINFORCE-based method, due to computational constraints. Through experiments on a curated small set of carefully selected prompts, we highlight GRPO’s weakness in overfitting specific prompts. Our subsequent experiments on long-form CoT tasks demonstrate REINFORCE++’s superior performance on challenging test datasets, indicating better Out-of-Distribution (OOD) generalization. In Section 5, we conclude our method and future work directions.
Figure 1: The comparison of PPO, Remax, GRPO, RLOO, and REINFORCE++.
2 Background and Related Work
State-of-the-art applications such as ChatGPT/GPT-4, Claude, and Gemini utilize reinforcement learning algorithms, such as PPO, for policy optimization. In particular, PPO optimizes LLMs by maximizing the following surrogate objective:
$\mathcal{L}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(O|q)}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}\min\left(s_{t}(\theta)A_{t},\ \text{clip}(s_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$ (1)
where:
$s_{t}(\theta)=\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}$ (2)
PPO is an actor-critic-based reinforcement learning (RL) algorithm where the critic model consumes significant training resources. Consequently, many researchers have proposed a series of REINFORCE-based methods, such as ReMax, RLOO, and GRPO, to avoid the computational overhead associated with the critic model while still obtaining relatively accurate token-wise advantage estimations. These methods design alternative techniques to calculate the baseline reward for each prompt as the advantage estimation.
Specifically, the advantage is computed to evaluate the performance of an action relative to a baseline in PPO. It starts by collecting samples of states, actions, rewards, and the following states. The advantage is then calculated using Generalized Advantage Estimation (GAE), which combines the temporal difference error $\delta_{q,o_{t}}=r_{t}+\gamma V(o_{t+1})-V(o_{t})$ over a series of time steps:
$A_{q,o_{t}}=\sum_{l=0}^{\infty}(\gamma\lambda)^{l}\delta_{t+l}$ (3)
where $\lambda$ is a parameter balancing bias and variance. To obtain an accurate advantage estimation and generalize to unseen tokens, PPO must use the critic model to learn the advantage function, obtaining an accurate advantage estimation and generalizing to unseen tokens.
As shown in Figure 1, ReMax adopts the greedy decoding method to generate a response and obtain its reward as the baseline reward for this prompt. Accordingly, the advantage of a query $q$ is given by
$A_{q,o_{t}}=r(o_{1:T},q)-r(\hat{o}_{1:T},q)$ (4)
where:
$\hat{o}_{1:T}=\operatorname*{argmax}_{o^{\prime}_{1},\dots,o^{\prime}_{T}}\prod_{i=1}^{T}\pi_{\theta}(o^{\prime}_{i}\mid q,o^{\prime}_{<i})$ (5)
Notably, ReMax does not use this response to train the model.
Both RLOO and GRPO sample multiple responses for a prompt. RLOO adopts the average rewards of all other samples for the current prompt as a baseline, which is shown in Equation 6.
$A_{q,o_{t}^{(i)}}=r(o_{1:T}^{(i)},q)-\frac{1}{k-1}\sum_{j\neq i}r(o_{1:T}^{(j)},q)$ (6)
GRPO, on the other hand, adopts the group relative advantage estimation method, which uses the mean reward divided by the standard deviation of all sampled responses for the current prompt as the baseline reward, which is shown in Equation 7.
$A_{q,o_{t}^{(i)}}=\frac{r(o_{1:T}^{(i)},q)-\text{mean}\bigl{(}\{r(o_{1:T}^{(j)},q)\}_{j=1}^{k}\bigr{)}}{\text{std}\bigl{(}\{r(o_{1:T}^{(j)},q)\}_{j=1}^{k}\bigr{)}}$ (7)
However, given a diverse dataset, we consider that the baseline reward for each prompt is not essential in the RLHF setting. Although the baseline reward of each prompt can provide relatively accurate advantage estimation for each training prompt and help the model learn the response with the highest reward under a prompt, it also accelerates the reward hacking and overfitting problems. It is worth noting that RLHF has two key differences from traditional RL problems:
  • In traditional RL problems, we train an RL policy and test it in the same environment. However, the RLHF method trains a prompt set and tests on another dataset, even an out-of-distribution (OOD) dataset.
  • In traditional RL problems, there is always a golden reward. In contrast, the RLHF method always uses a reward model or rule-based reward, which may encounter reward hacking.
Accordingly, both reward hacking and overfitting problems may degrade model performance. Specifically, optimizing multiple responses for a prompt in a batch using methods like RLOO and GRPO tends to overfit the best response under specific simple prompts, ultimately degrading the model’s generalization. Furthermore, optimizing multiple responses for a prompt within a single batch reduces the diversity of the model’s outputs, leading to a low diversity in token-level advantage distribution and causing overfitting on those tokens. In contrast, PPO does not suffer significantly from this issue, as the value network continues to be trained and retains the learned advantages, thereby supporting a more generalizable token-wise advantage.
Accordingly, to avoid overfitting specific prompts and increase prompt diversity in a training batch, REINFORCE++ can sample one response for each prompt and normalize the token-wise advantage in the global batch size to increase training stability.
3 Method
REINFORCE++ optimizes the PPO objective and employs the clipping strategy defined in Equation 1. To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward. Accordingly, the advantage of REINFORCE++ is the normalized reward-to-go in reinforcement learning:
$A_{q,o_{t}}=r(o_{1:T},q)-\beta\cdot\sum_{i=t}^{T}\text{KL}(i)$ (8)
where:
$\text{KL}(t)=\log\left(\frac{\pi_{\theta_{\text{old}}}^{\text{RL}}(o_{t}|q,o_{<t})}{\pi^{\text{SFT}}(o_{t}|q,o_{<t})}\right)$ (9)
The gradient of the token-level KL penalty has been theoretically proven to be unbiased concerning the $k_{3}$ loss of GRPO in RLHF. Additionally, we normalize this advantage across the global batch for all prompts:
$A_{q,o_{t}}^{\text{norm}}=\frac{A_{q,o_{t}}-\text{mean}\left(A_{q,o_{t}}\right)}{\text{std}\left(A_{q,o_{t}}\right)},$ (10)
REINFORCE++ removes the prompt-specific reward baseline and does not optimize multiple responses from a single prompt within a training batch. Compared to PPO, REINFORCE++ eliminates the critic model and sets the discount factor to $1$. The detailed implementation of our algorithm is provided in Algorithm 1, with further algorithmic details discussed therein.
Algorithm 1 REINFORCE++
1:Initial policy model $\pi_{\text{init}}$, reward models $R$, task prompts $\mathcal{D}$
2:policy model $\pi_{\theta}\leftarrow\pi_{\text{init}}$
3:for step $=1,\ldots,M$ do
4:     Sample a batch $\mathcal{D}_{b}$ from $\mathcal{D}$
5:     Update the old policy model $\pi_{\text{old}}\leftarrow\pi_{\theta}$
6:     Sample one output $o\sim\pi_{\text{old}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{b}$
7:     Compute rewards $r_{i}$ for each sampled in a batch of output $o_{i}$ by running $R$
8:     Compute $A_{q,o_{t}}^{norm}$ for the $t$-th token of $o$ for prompt $q$ through Equation 8 to 10
9:     for iteration $=1,\ldots,k$ do
10:         Update the policy model $\pi_{\theta}$ by maximizing the REINFORCE++ objective (Equation 1)
11:     end for
12:end for
13:$\pi_{\theta}$
4 Experiments
The empirical evaluation of REINFORCE++ was conducted across diverse test scenarios to assess its performance comprehensively. In experiments using the Bradley-Terry Reward Model, we thoroughly compare REINFORCE++ against existing methods, including ReMax, RLOO, GRPO, and PPO. Due to computational constraints, we focused our comparisons between REINFORCE++ and GRPO, the current state-of-the-art REINFORCE-based method for experiments in the long-form CoT setting.
4.1 Performance with Outcome Reward Model
Experimental Setup
This study investigates reinforcement learning with human feedback (RLHF) in a general-domain setting. The experimental process began with an instruction-following policy model, fine-tuned on diverse datasets reflecting general-domain language tasks. The model was subsequently refined using reinforcement learning algorithms guided by a reward model. Specifically, a Bradley-Terry reward model was trained on human-generated preference data obtained through pairwise comparisons of model outputs. The reward model evaluates the policy model’s responses based on helpfulness, accuracy, coherence, and alignment with human intent and outputs an outcome reward value for each Bai et al. (2022).
Reward Model
Following the approach proposed by Ouyang et al. (2022b), we initialize the reward model using a SFT model  [^1] https://huggingface.co/OpenRLHF/Llama-3-8b-sft-mixture. To adapt the model for preference learning, we replace the final layer with a linear head that produces a scalar output. The reward model is then trained using the negative log-likelihood loss function defined as:
$\mathcal{L}_{\text{RM}}(\theta)=-\mathbb{E}_{\left(q,o^{+},o^{-}\right)\sim\mathcal{D}}\left[\log\sigma\left(r_{\theta}\left(q,o^{+}\right)-r_{\theta}\left(q,o^{-}\right)\right)\right]$ (11)
Dataset
For the training of our reward model, we leveraged an extensive dataset comprising approximately 700,000 pairs of human preference data [^2] https://huggingface.co/datasets/hendrydong/preference_700K aggregated from multiple publicly available datasets. These datasets provide a rich diversity of contexts and preferences, facilitating the reward model in effectively capturing nuanced human judgment. To systematically prompt the policy model for response generation, we curated a carefully balanced set of 20,000 prompts sampled from diverse sources [^3] https://huggingface.co/datasets/RLHFlow/prompt-collection-v0.1, ensuring comprehensive coverage of various scenarios and domains. This diversity promotes robustness and generalization in the responses generated by the policy model.
Score Length Per Token Score
REINFORCE++ 46.7 832 0.0561
REINFORCE++ Baseline 44.2 834 0.0530
GRPO 46.8 860 0.0544
RLOO 44.6 866 0.0515
ReMax 45.1 805 0.0560
Table 1: Comparison between GRPO and REINFORCE++ on Score and Length. Better results for each reward model are highlighted in bold.
Experimental Results
We use Chat-Arena-Hard Li et al. (2024) to evaluate our models. As illustrated in Table 1, GRPO achieves a slightly superior overall score of 46.8 compared to 46.7 for REINFORCE++. However, GRPO generates longer sequences of an average of 860 tokens, whereas REINFORCE++ produces shorter outputs of only 832 tokens. Consequently, when evaluating performance on a per-token basis, REINFORCE++ outperforms GRPO, achieving a higher per-token score of 0.0561 compared to GRPO’s 0.0544. The table suggests that REINFORCE++ provides more efficient outputs despite the preference bias of the evaluation model on the length.
Figure 2: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and KL Divergence with Outcome Reward Model.
Results Analysis
We plot the test reward and KL divergence curves in Figure 2 to provide a clear insight into the comparative behavior of GRPO and REINFORCE++. Initially, GRPO achieves significantly higher rewards and outperforms REINFORCE++ throughout training. However, closer inspection reveals that the superior rewards achieved by GRPO are mainly due to reward hacking. Specifically, the rapid increase in KL divergence of GRPO suggests that the model is hacking the reward signal rather than improving its generalization performance, resulting in inflated reward values without corresponding gains on the test set. In contrast, REINFORCE++ shows a more gradual but stable increase in reward, accompanied by a modest rise in KL divergence. The result indicates a favorable trade-off, with substantial improvements in the reward being achieved with minimal deviation from the reference model. The comparatively small increase in KL divergence for REINFORCE++ highlights that each divergence unit is translated into a more effective and robust policy improvement, suggesting a higher KL-to-reward conversion efficiency. To substantiate our argument, we evaluated our model on Out-of-Distribution tasks with the setting of RLHF, which primarily includes mathematical problem tasks (GSM8K, MATH) and code generation tasks (HumanEval, MBPP). The evaluation results are presented in Table 2.
GSM8K MATH HumanEval MBPP Avg.
REINFORCE++ 96.21 75.20 85.98 84.39 85.45
REINFORCE++ Baseline 95.98 72.40 78.66 85.45 83.12
GRPO 96.21 73.80 80.49 83.33 83.46
RLOO 96.44 72.40 79.27 82.54 82.67
ReMax 96.59 75.40 78.66 82.54 84.05
Table 2: Comparison between different advantage estimate methods on OOD benchmarks. Better results for each reward model are highlighted in bold.
4.2 Performance on Long-form Chain-of-Thought Tasks
4.2.1 Analysis on Small-Scale Datasets
Experiment Setup
We trained the Qwen2.5-Math-7B pre-trained model using only 30 questions and answers from AIME-24 and evaluated its performance on the AIME-25 dataset. Under this limited training setting, GRPO notably demonstrated its weakness by overfitting the small training dataset.
AIME-24 AIME-25
Pass@N N = 1 N = 1 N = 16
GRPO 95.0 0.0 0.4
REINFORCE++ 71.0 2.5 40.0
Table 3: Comparison between GRPO and REINFORCE++ on both training and test datasets (AIME-24 and AIME-25). Better results for each reward model are highlighted in bold.
Experimental Results
As shown in Table 3, GRPO achieves nearly perfect scores (approximately 100) on the training dataset (AIME-24). However, it performs poorly on the testing dataset (AIME-25), scoring almost 0 in both Pass@1 and Pass@16 test settings. In contrast, while REINFORCE++ achieves a more modest score of 71.0 on AIME-24, it demonstrates better generalization with scores of 2.5 and 40.0 on Pass@1 and Pass@16 test settings, respectively.
Figure 3: Training curves for 15 randomly selected questions during reinforcement learning from zero on a small prompt dataset. Left: Scores trained using GRPO; Right: Scores trained using REINFORCE++
Results Analysis
Furthermore, we analyze the training curves of 15 randomly selected cases. Our findings reveal that GRPO rapidly achieves 100% accuracy (represented by a 1.0 score in Figure 3) within a few steps. In contrast, REINFORCE++ demonstrates a more gradual improvement, typically reaching the same level of accuracy throughout 10 to 20 steps. These observations suggest that GRPO overfits the training set. Further analysis reveals that responses from the GRPO-trained model are significantly shorter, averaging only 30 tokens, compared to 425 tokens for the REINFORCE++-trained model. This evidence confirms that GRPO is more susceptible to overfitting training prompts than REINFORCE++, resulting in poorer generalization performance on test datasets.
4.2.2 RL from Supervised Fine-tuned Model
Experimental Setup
In real-world tasks, user prompts exhibit enormous form and intent, making it difficult to assess a given model’s strengths and weaknesses systematically. Using synthetic datasets, we can systematically manipulate key factors such as length and difficulty, allowing for a more direct and interpretable evaluation and analysis of model performance.
Dataset & Hyper-Parameter
Following Logic-RL (Xie et al., 2025), we incorporate the Knights and Knaves (K&K) puzzles (Xie et al., 2024) into RL training as an algorithmically generated dataset for logical reasoning. In these puzzles, each character is either a knight, who always tells the truth, or a knave, who always lies. The objective is to determine each character’s identity based on their statements. A key feature of this dataset is its strong controllability. The length of the prompt is proportional to the number of roles, and the difficulty can be adjusted by modifying the complexity of logical operations. These puzzles serve as unseen data for the original model, making them ideal for evaluating generalization capabilities. Since the model’s performance in logical reasoning depends on its ability to follow instructions and understand context, we did not begin our experiment with the base model. Instead, we selected a model, Qwen2.5-7B-Instruct-1M, with enhanced capability for handling more extended contexts and following control instructions. We keep all hyper-parameters the same and compare reinforcement learning algorithms by varying the advantage estimation function.
Figure 4: Comparison between GRPO and REINFORCE++ on logic benchmarks with different difficulty levels. The lines show the average performance of the two methods.
Experimental Results
Figure 4 presents a comparative analysis of GRPO and REINFORCE++ on the test datasets. Increasing the number of people - a proxy for task difficulty - leads to decreased performance for both methods. In particular, GRPO achieves slightly higher scores in simpler scenarios involving two or three people, but its performance deteriorates significantly as the number of people increases. On the other hand, REINFORCE++ shows more excellent performance stability and achieves better results in scenarios involving four or more people. This trend is particularly evident in the Out-of-Distribution scenario of eight people, a setting not present in the training data, where GRPO scores 20 and REINFORCE++ scores 36. Overall, REINFORCE++ achieves an average score of 62.1, outperforming the average of 55.7. These results indicate that REINFORCE++ generalizes more effectively in complex and OOD scenarios, underscoring its superior robustness compared to GRPO.
Results Analysis
We analyze the training curves in Figure 5. Our findings indicate that GRPO achieves significant reward values within a few hundred steps. In contrast, REINFORCE++ shows a more moderate and gradual increase in reward, ultimately converging to a higher stable value. Regarding response length, the GRPO-trained model generates considerably shorter responses, averaging approximately 600 tokens, whereas REINFORCE++ consistently produces longer responses of around 1000 tokens. This disparity suggests that GRPO generates shorter, potentially superficial responses, hinting at possible overfitting or memorization rather than meaningful reasoning. Consequently, the model trained with GRPO might generalize poorly on unseen test cases, reinforcing our earlier observations of its susceptibility to overfitting.
4.2.3 RL from Zero Setting
Experimental Setup
Inspired by DeepSeek-R1 (Guo et al., 2025), we apply reinforcement learning with verified reward (RLVR) to a base model for mathematical tasks to evaluate different strategies on the model’s reasoning ability. Previous studies suggest starting from a powerful base model with strong reasoning potential; therefore, we choose Qwen2.5-Math-Base [^4] https://huggingface.co/Qwen/Qwen2.5-Math-7B as the base model.
Figure 5: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and Response Length from Supervised Fine-tuned Model. The $x$ axis represents the training steps.
Dataset
The dataset primarily comprises difficulty levels 3 to 5 of the MATH training split. Due to the context length limitations of the base model, we select approximately 8,000 shorter prompts from the dataset to maximize the observation of variations in the model’s output length.
Hyper-Parameter
We keep all hyper-parameters the same and compare reinforcement learning algorithms by varying the advantage estimation function. In each exploration step, we select 32 questions and generate eight answers per question to maintain a balance between computational efficiency and stability. The actor learning rate is $5\times 10^{-7}$ to optimize the convergence of policy updates. A discount factor ($\gamma$) 1.0 is used to emphasize long-term reward accumulation. The clipping parameter ($\epsilon$) is set to 0.2 to stabilize training and regulate policy updates. The KL penalty coefficient ($\beta$) is set to 0.001 to ensure controlled divergence from the reference model.
AIME-24 AMC-23 MATH-500
Pass@N N = 8 N = 8 N = 1
GRPO 18.96 59.22 73.00
REINFORCE++ 21.04 60.47 72.00
Table 4: Comparison between GRPO and REINFORCE++ on both training and test datasets (AIME-24 and AIME-25). Better results for each reward model are highlighted in bold.
Experimental Results & Analysis
As shown in table 4, REINFORCE++ consistently outperforms GRPO in the out-of-distribution setting and performs similarly in the other setting. Specifically, on the in-distribution test dataset (MATH-500), GRPO achieves a slightly higher Pass@1 score of 73.00 compared to 72.00 for REINFORCE++. However, REINFORCE++ demonstrates superior generalization on the OOD scores, outperforming GRPO on AIME-24 with a Pass@8 score of 21.04 compared to 18.96 and on AIME-25 with a score of 60.47 compared to 59.22. These results indicate that GRPO performs competitively within the training distribution but shows signs of overfitting, whereas REINFORCE++ generalizes better to novel, particularly challenging OOD scenarios.
Figure 6: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and Response Length with Rule Reward Model. The $x$ axis represents the training steps.
5 Conclusion
This paper presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences. Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global batch mean reward as a baseline to prevent overfitting and instability. Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities. Future work includes exploring adaptive normalization techniques, advanced variance reduction methods, and extending REINFORCE++ beyond RLHF settings.
Appendix A Algorithm Details
REINFORCE++-baseline.
For ease of use and reproducibility, we implement our proposed method within the OpenRLHF framework (https://github.com/OpenRLHF/OpenRLHF).
Our experiments above showed that training models using multiple generated responses per prompt can further improve performance. Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation. Specifically, we sample multiple responses for each prompt and compute their average reward as the baseline to reshape the rewards. The advantage for each prompt-response pair is then defined as:
$Adv_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$
$Adv_{q,o_{t}}^{norm}=\frac{Adv_{q,o_{t}}-\text{mean}_{batch}(Adv_{q,o_{t}})}{\text{std}_{batch}(Adv_{q,o_{t}})}$
where $group$ denotes the generated responses corresponding to the same prompt. This baseline calculation is similar to the approach taken by GRPO; however, we deliberately move the standard deviation ($std$) of the group normalization to the global batch normalization. In practice, we observed that including the standard deviation introduces training instability and exacerbates overfitting to easier prompts in the RLHF process. For the KL constraint, REINFORCE++-baseline uses $k_{2}$ instead of the $k_{3}$ divergence approximation method in GRPO. For more details, please refer to the next paragraph.
KL Penalty Design.
In the design of REINFORCE-based RLHF methods, the choice of KL divergence estimation significantly influences algorithmic stability and accuracy. The widely utilized $k_{3}$ estimation in GRPO, defined by the loss function:
$\mathcal{L}_{k_{3}}=\mathbb{E}_{x\sim D,y\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left(\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-\log\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-1\right)$ (12)
Which involve inherent biases due to its approximation nature. Specifically, the gradient of $k_{3}$ estimation approximates the gradient of the $k_{2}$ estimation through linearization:
$k_{2}$ Gradient $:\log x\cdot\nabla_{\theta}\log\pi_{\theta},$ (13)
$k_{3}$ Gradient $:(x-1)\cdot\nabla_{\theta}\log\pi_{\theta},$ (14)
where $x=\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$. Although the $k_{3}$ gradient aligns closely with the $k_{2}$ gradient when the policy is close to the reference ($x\approx 1$), substantial deviations result in two key deficiencies:
  1. 1.
    Bias: As the current policy diverges significantly from the reference, particularly when $\pi_{\text{ref}}\gg\pi_{\theta_{\text{old}}}$, the linear approximation error expands nonlinearly, causing considerable biases.
  2. 2.
    Asymmetry: The function $(x-1)$ exhibits asymmetric behavior depending on whether $\pi_{\theta_{\text{old}}}$ surpasses or falls below $\pi_{\text{ref}}$, further complicating convergence properties.
Therefore, despite its computational simplicity, $k_{3}$ estimation introduces biases and increased variance, suggesting it is not strictly superior to the theoretically unbiased $k_{2}$estimation. Experimental evidence consistently reveals that employing $k_{3}$ estimation in GRPO results in higher variance fluctuations than $k_{2}$ estimation. Further discussion can be found in Liu (2025).
Relationship with PPO.
REINFORCE++ shares certain similarities with PPO in its formulation. Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and employs global batch normalization as the baseline. Mathematically, this relationship is expressed as follows:
$\text{GAE}(\lambda=1,\gamma=1)=\sum_{l=0}^{\infty}r_{t+l}-V(s_{t}),$ (15)
where $V(s_{t})$ denotes the estimated value function at state $s_{t}$. By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term. At the same time, the introduction of global batch normalization further stabilizes training by reducing variance and preventing overfitting to individual prompts.
Other Implementation Details
Mini-Batch Updates
To enhance training efficiency, we implement mini-batch updates with the following characteristics:
  • Batch Processing: Data is processed in smaller, manageable chunks rather than full-batch updates.
  • Multiple Updates: Each mini-batch allows for multiple parameter updates, improving convergence rates.
  • Stochastic Optimization: Introduces beneficial randomness for better generalization.
Reward Normalization and Clipping
We implement comprehensive reward processing to stabilize training:
  • Normalization: Standardizes rewards using z-score normalization to mitigate outliers.
  • Clipping: Constrains reward values within predefined bounds to avoid instability.
  • Scaling: Applies appropriate scaling factors for numerical stability during updates.
Appendix B Acknowledgements
  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the early draft of the paper.
  • Jason Klein Liu: Implemented the experimental code, fine-tuned hyperparameters, wrote the paper, and provided GPU resources.
  • Wei Shen: Led the paper writing and supervised the main experiments.

### v8

@@
 Abstract
-Reinforcement Learning from Human Feedback (RLHF) is crucial in aligning large language models (LLMs) with human values and preferences.
-While state-of-the-art applications like ChatGPT/GPT-4 commonly employ Proximal Policy Optimization (PPO), including a critic network introduces significant computational overhead.
+Reinforcement Learning from Human Feedback (RLHF) plays a crucial role in aligning large language models (LLMs) with human values and preferences.
+While state-of-the-art applications like ChatGPT or GPT-4 commonly employ Proximal Policy Optimization (PPO), the inclusion of a critic network introduces significant computational overhead.
 REINFORCE-based methods, such as REINFORCE Leave One-Out (RLOO), ReMax, and Group Relative Policy Optimization (GRPO), address this limitation by eliminating the critic network.
 However, these approaches face challenges in accurate advantage estimation.
-Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on more straightforward prompts and vulnerability to reward hacking.
-To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the normalized reward of a batch as the baseline.
+Specifically, they estimate advantages independently for responses to each prompt, which can lead to overfitting on simpler prompts and vulnerability to reward hacking and may be biased.
+To address these challenges, we introduce REINFORCE++, a novel approach that removes the critic model while using the global advantage normalization which is unbiased to improve the training stability.
 Our empirical evaluation demonstrates that REINFORCE++ exhibits robust performance across various reward models without requiring prompt set truncation.
-Furthermore, it achieves superior generalization in RLHF and long chain-of-thought (CoT) settings compared to REINFORCE-based methods.
+Furthermore, it achieves superior generalization in both RLHF and long chain-of-thought (CoT) settings compared to existing REINFORCE-based methods.
 The implementation is available at https://github.com/OpenRLHF/OpenRLHF.
 1 Introduction
-Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning large language models (LLMs) with human values and preferences  (Vemprala et al., 2023; Achiam et al., 2023; Ouyang et al., 2022a; Shen & Zhang, 2024).
+Reinforcement Learning from Human Feedback (RLHF) is a key technique for aligning large language models (LLMs) with human values and preferences  (Vemprala et al., 2023; Achiam et al., 2023; Ouyang et al., 2022a; Shen & Zhang, 2024; Shen et al., 2025; Hu et al., 2024).
 Despite the emergence of non-RL alternatives like DPO  (Rafailov et al., 2023), state-of-the-art applications such as ChatGPT/GPT-4  (Vemprala et al., 2023; OpenAI, 2023), Claude  (Anthropic, 2023), and Gemini  (Team et al., 2023) continue to rely on RL algorithms, particularly PPO, for policy optimization.
 However, PPO  (Schulman et al., 2017) requires a critic network, introducing substantial computational overhead and memory demands that limit large model alignment in small-scale clusters.
-To address this, researchers have proposed various REINFORCE-based methods that eliminate the critic network, including ReMax  (Li et al., 2023), REINFORCE Leave One-Out (RLOO)  (Wu et al., 2024), and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024).
-Furthermore, Deepseek-R1 demonstrates the effectiveness of the REINFORCE-based method in the long-form CoT setting  (Guo et al., 2025), which achieved state-of-the-art performance on challenging datasets using GRPO with rule-based rewards.
+To address this, researchers have proposed various REINFORCE-based methods that eliminate the critic network, including ReMax  (Li et al., 2023), REINFORCE Leave One-Out (RLOO)  (Ahmadian et al., 2024), and Group Relative Policy Optimization (GRPO)  (Shao et al., 2024).
+Furthermore, Deepseek-R1 demonstrates the effectiveness of the REINFORCE-based method in the long-form CoT setting  (Guo et al., 2025; Seed et al., 2025), which achieved state-of-the-art performance on challenging datasets using GRPO with rule-based rewards.
 Without the critic network, REINFORCE-based methods often struggle to estimate individual tokens’ advantages accurately.
-Various REINFORCE-baseline approaches have been proposed to address this limitation, but each has significant drawbacks.
+Various REINFORCE-baseline approaches have been proposed to address this limitation, among which are those that introduce prompt-level baselines; yet each still has significant drawbacks.
 ReMax uses a greedy search to generate a response for each prompt and employs its reward as the baseline, inefficiently consuming a model response solely for baseline computation.
-RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO utilizes a normalized reward across all responses.
+RLOO and GRPO take a different approach by generating multiple responses per prompt: RLOO uses the mean reward of other responses as the baseline, while GRPO employs a normalized reward across all responses, this serves as a biased estimator of the advantage (see Appendix A).
 These methods improve advantage estimation accuracy, but their practice of optimizing multiple responses per prompt intensifies the risk of reward hacking.
@@
 Therefore, these methods require carefully curating prompt sets particular to each task.
-To address these challenges, we propose REINFORCE++, a novel REINFORCE-based method that eliminates the critic model from PPO and uses the mean reward of a global batch as the baseline.
-This approach prevents overfitting to specific training prompts and demonstrates robustness across both Bradley-Terry and rule-based reward models.
+To address these challenges, we propose REINFORCE++, a novel REINFORCE-based method that eliminates the critic model from PPO and uses the global advantage normalization.
+This approach is unbiased and prevents overfitting to specific training prompts and demonstrates robustness across both Bradley-Terry and rule-based reward models.
 Notably, REINFORCE++ eliminates the need for prompt set truncation and achieves strong generalization performance in both RLHF and long CoT RL settings.
-The remainder of this paper is organized as follows.
-In Section 2, we analyze the differences between various REINFORCE-based RLHF methods, demonstrating the ineffectiveness of prompt-specific reward baselines and the overfitting problems in RLOO and GRPO methods.
-Section 3 presents REINFORCE++, detailing its implementation and advantages.
-In Section 4, we evaluate REINFORCE++ through comprehensive experiments.
-We first compare REINFORCE++ with other RLHF methods using both Bradley-Terry and Rule-Based Reward Models, showing its superior or comparable performance.
-Then, under the long-form Chain-of-Thought (CoT) setting, we focus on comparing REINFORCE++ with GRPO, the current state-of-the-art REINFORCE-based method, due to computational constraints.
-Through experiments on a curated small set of carefully selected prompts, we highlight GRPO’s weakness in overfitting specific prompts.
-Our subsequent experiments on long-form CoT tasks demonstrate REINFORCE++’s superior performance on challenging test datasets, indicating better Out-of-Distribution (OOD) generalization.
-In Section 5, we conclude our method and future work directions.
+In summary, our contributions are as follows:
+  • We analyze the limitations of existing REINFORCE-based RLHF methods, revealing that prompt-specific reward baselines are ineffective and identifying overfitting issues in RLOO and GRPO.
+  • We propose REINFORCE++, a novel REINFORCE-based RLHF method, and detail its implementation, highlighting its advantages in addressing the identified limitations.
+  • Through comprehensive experiments using both Bradley-Terry and Rule-Based Reward Models, we demonstrate that REINFORCE++ achieves superior or comparable performance compared to other RLHF methods.
+  • In long-form Chain-of-Thought (CoT) settings, we show that GRPO suffers from overfitting to specific prompts, while REINFORCE++ exhibits better Out-of-Distribution (OOD) generalization, especially on challenging test datasets.
+2 Background and Related Work
 Figure 1: The comparison of PPO, Remax, GRPO, RLOO, and REINFORCE++.
-2 Background and Related Work
-State-of-the-art applications such as ChatGPT/GPT-4, Claude, and Gemini utilize reinforcement learning algorithms, such as PPO, for policy optimization.
+State-of-the-art applications such as ChatGPT or GPT-4, Claude, and Gemini utilize reinforcement learning algorithms, such as PPO, for policy optimization.
 In particular, PPO optimizes LLMs by maximizing the following surrogate objective:
-$\mathcal{L}_{\text{PPO}}(\theta)=\mathbb{E}_{q\sim P(Q),o\sim\pi_{\theta_{\text{old}}}(O|q)}\left[\frac{1}{|o|}\sum_{t=1}^{|o|}\min\left(s_{t}(\theta)A_{t},\ \text{clip}(s_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$ (1)
+$\mathcal{L}_{\text{PPO}}(\theta)$ $=\mathbb{E}_{q\sim P(Q),\ o\sim\pi_{\theta_{\text{old}}}(O|q)}\Bigg{[}\frac{1}{|o|}\sum_{t=1}^{|o|}\min\Big{(}s_{t}(\theta)A_{t},$
+$\text{clip}(s_{t}(\theta),\ 1-\epsilon,\ 1+\epsilon)A_{t}\Big{)}\Bigg{]}$ (1)
 where:
 $s_{t}(\theta)=\frac{\pi_{\theta}(o_{t}|q,o_{<t})}{\pi_{\theta_{\text{old}}}(o_{t}|q,o_{<t})}$ (2)
-PPO is an actor-critic-based reinforcement learning (RL) algorithm where the critic model consumes significant training resources.
+PPO is an actor-critic-based reinforcement learning algorithm where the critic model consumes significant training resources.
 Consequently, many researchers have proposed a series of REINFORCE-based methods, such as ReMax, RLOO, and GRPO, to avoid the computational overhead associated with the critic model while still obtaining relatively accurate token-wise advantage estimations.
 These methods design alternative techniques to calculate the baseline reward for each prompt as the advantage estimation.
-Specifically, the advantage is computed to evaluate the performance of an action relative to a baseline in PPO.
+Specifically, PPO computes the advantage to assess how an action’s return compares to the immediately preceding state value.
 It starts by collecting samples of states, actions, rewards, and the following states.
@@
 where $\lambda$ is a parameter balancing bias and variance.
-To obtain an accurate advantage estimation and generalize to unseen tokens, PPO must use the critic model to learn the advantage function, obtaining an accurate advantage estimation and generalizing to unseen tokens.
+To obtain an accurate advantage estimation and generalize to unseen tokens, PPO use the critic model to learn the advantage function, obtaining an accurate advantage estimation and generalizing to unseen tokens.
 As shown in Figure 1, ReMax adopts the greedy decoding method to generate a response and obtain its reward as the baseline reward for this prompt. Accordingly, the advantage of a query $q$ is given by
@@
 Notably, ReMax does not use this response to train the model.
-Both RLOO and GRPO sample multiple responses for a prompt. RLOO adopts the average rewards of all other samples for the current prompt as a baseline, which is shown in Equation 6.
+In addition, both RLOO and GRPO sample multiple responses for a prompt.
+RLOO adopts the average rewards of all other samples for the current prompt as a baseline, which is shown in Equation 6.
 $A_{q,o_{t}^{(i)}}=r(o_{1:T}^{(i)},q)-\frac{1}{k-1}\sum_{j\neq i}r(o_{1:T}^{(j)},q)$ (6)
@@
 $A_{q,o_{t}^{(i)}}=\frac{r(o_{1:T}^{(i)},q)-\text{mean}\bigl{(}\{r(o_{1:T}^{(j)},q)\}_{j=1}^{k}\bigr{)}}{\text{std}\bigl{(}\{r(o_{1:T}^{(j)},q)\}_{j=1}^{k}\bigr{)}}$ (7)
-However, given a diverse dataset, we consider that the baseline reward for each prompt is not essential in the RLHF setting.
-Although the baseline reward of each prompt can provide relatively accurate advantage estimation for each training prompt and help the model learn the response with the highest reward under a prompt, it also accelerates the reward hacking and overfitting problems.
+For GRPO, which has been widely adopted in various works, we demonstrate that its advantage estimation is biased (see Appendix A for details).
+Furthermore, when working with a diverse dataset, we argue that the per-prompt baseline reward is not essential in the RLHF framework.
+While such a baseline can yield relatively accurate advantage estimates for each training prompt—thereby helping the model learn to generate responses with the highest reward under a given prompt—it also exacerbates reward hacking and overfitting issues.
 It is worth noting that RLHF has two key differences from traditional RL problems:
@@
 Specifically, optimizing multiple responses for a prompt in a batch using methods like RLOO and GRPO tends to overfit the best response under specific simple prompts, ultimately degrading the model’s generalization.
-Furthermore, optimizing multiple responses for a prompt within a single batch reduces the diversity of the model’s outputs, leading to a low diversity in token-level advantage distribution and causing overfitting on those tokens.
+Furthermore, when optimizing multiple responses for a single prompt within a training batch, the diversity of the model’s outputs during training is reduced.
+This leads to low diversity in the token-level advantage distribution, ultimately causing overfitting on those tokens.
 In contrast, PPO does not suffer significantly from this issue, as the value network continues to be trained and retains the learned advantages, thereby supporting a more generalizable token-wise advantage.
@@
 3 Method
-REINFORCE++ optimizes the PPO objective and employs the clipping strategy defined in Equation 1.
+REINFORCE++ still optimizes the PPO objective and employs the clipping strategy as defined in Equation 1.
 To further reduce the variance of the gradient estimate, making the learning process more stable and efficient, we adopt the average reward of a global training batch as the baseline reward.
@@
 $\text{KL}(t)=\log\left(\frac{\pi_{\theta_{\text{old}}}^{\text{RL}}(o_{t}|q,o_{<t})}{\pi^{\text{SFT}}(o_{t}|q,o_{<t})}\right)$ (9)
-The gradient of the token-level KL penalty has been theoretically proven to be unbiased concerning the $k_{3}$ loss of GRPO in RLHF.
+Notably, Reinforce++ algorithm employs the KL-based $k_{1}$ loss.
+This choice is motivated by the fact that the GRPO algorithm, which relies on the KL-based $k_{3}$ loss, suffers from bias in its gradient estimation (For a detailed analysis, refer to Appendix KL Penalty Design.).
 Additionally, we normalize this advantage across the global batch for all prompts:
-$A_{q,o_{t}}^{\text{norm}}=\frac{A_{q,o_{t}}-\text{mean}\left(A_{q,o_{t}}\right)}{\text{std}\left(A_{q,o_{t}}\right)},$ (10)
-REINFORCE++ removes the prompt-specific reward baseline and does not optimize multiple responses from a single prompt within a training batch.
-Compared to PPO, REINFORCE++ eliminates the critic model and sets the discount factor to $1$.
-The detailed implementation of our algorithm is provided in Algorithm 1, with further algorithmic details discussed therein.
+$A_{q,o_{t}}^{\text{norm}}=\frac{A_{q,o_{t}}-\text{mean}\left(A_{q,o_{t}}\mid A_{q,o_{t}}\in\mathcal{D}_{\text{batch}}\right)}{\text{std}\left(A_{q,o_{t}}\mid A_{q,o_{t}}\in\mathcal{D}_{\text{batch}}\right)}$ (10)
+The global advantage normalization helps avoid training instability caused by excessively large advantage values (Andrychowicz et al., 2020).
+Since the global batch size is typically large, we can consider the mean and variance as constants without introducing bias into the policy gradient estimation.
+For further analysis, see the Appendix A.2.
+It is worth noting that, compared to PPO, REINFORCE++ essentially eliminates the critic model and sets the GAE discount factor to 111.
+The detailed implementation of our algorithm is provided in Algorithm 1, with further algorithmic details discussed in Appendix Algorithm Details.
 Algorithm 1 REINFORCE++
@@
 3:for step $=1,\ldots,M$ do
-4:     Sample a batch $\mathcal{D}_{b}$ from $\mathcal{D}$
+4:  Sample a batch $\mathcal{D}_{batch}$ from $\mathcal{D}$
 5:  Update the old policy model $\pi_{\text{old}}\leftarrow\pi_{\theta}$
-6:     Sample one output $o\sim\pi_{\text{old}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{b}$
+6:  Sample one output $o\sim\pi_{\text{old}}(\cdot\mid q)$ for each question $q\in\mathcal{D}_{batch}$
 7:  Compute rewards $r_{i}$ for each sampled in a batch of output $o_{i}$ by running $R$
@@
 12:end for
-13:$\pi_{\theta}$
+[…]
+3.1 REINFORCE++-Baseline
+Recent research (Yue et al., 2025) showed that training models using multiple generated responses per prompt in reasoning tasks can further improve performance.
+Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation.
+Specifically, for each prompt we sample multiple responses and use their average reward as a baseline to reshape the rewards, therefore, under the RLVR setting, the reward function is not very sensitive whether it is 0/1 or -1/1.
+The advantage value for each prompt-response pair is then defined as:
+$A_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$ (11)
+$A_{q,o_{t}}^{norm}=\frac{A_{q,o_{t}}-\text{mean}_{batch}(A_{q,o_{t}})}{\text{std}_{batch}(A_{q,o_{t}})}$ (12)
+where $group$ denotes the generated responses corresponding to the same prompt.
+This baseline calculation is similar to the approach taken by GRPO; However, instead of using the local standard deviation in GRPO, we adopt the global batch normalization from REINFORCE++ to enhance training stability.
+For the REINFORCE++-baseline, we adopt the k2 KL estimator rather than k3 in GRPO, as k2 provides an unbiased estimate.
+See Appendix KL Penalty Design for more details.
+3.2 Relationship with PPO
+REINFORCE++ shares certain similarities with PPO in its formulation.
+Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and additionally employs global batch normalization as the baseline.
+Mathematically, this relationship is expressed as follows:
+$\text{GAE}(\lambda=1,\gamma=1)=\sum_{l=0}^{\infty}r_{t+l}-V(s_{t}),$ (13)
+where $V(s_{t})$ denotes the estimated value function at state $s_{t}$.
+By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term, while the introduction of global batch normalization further stabilizes training.
 4 Experiments
-The empirical evaluation of REINFORCE++ was conducted across diverse test scenarios to assess its performance comprehensively.
-In experiments using the Bradley-Terry Reward Model, we thoroughly compare REINFORCE++ against existing methods, including ReMax, RLOO, GRPO, and PPO.
-Due to computational constraints, we focused our comparisons between REINFORCE++ and GRPO, the current state-of-the-art REINFORCE-based method for experiments in the long-form CoT setting.
+The empirical evaluation of REINFORCE++ was conducted across diverse test scenarios to comprehensively assess its performance.
+In the context of RLHF, experiments utilizing the Bradley-Terry Reward Model were designed to thoroughly compare REINFORCE++ against existing methods, including ReMax, RLOO, GRPO, and PPO.
+For RLVR settings—specifically in long-form CoT experiments—our comparisons focused on REINFORCE++ versus GRPO, the current state-of-the-art REINFORCE-based method in this domain, due to computational constraints.
+We utilize binary (-1/1) rewards as training signals for the long-form CoT experiments.
+Additionally, we provide comparative experiments for our method under complex multi-turn tool call scenarios to demonstrate strong performance in agentic environments.
+This multi-faceted approach ensures a robust evaluation of REINFORCE++ and its variant across various reinforcement learning paradigms and application contexts.
+We outline the recommended best practices for our methods in Appendix D.
 4.1 Performance with Outcome Reward Model
 Experimental Setup
-This study investigates reinforcement learning with human feedback (RLHF) in a general-domain setting.
+This study investigates RLHF in a general-domain setting.
 The experimental process began with an instruction-following policy model, fine-tuned on diverse datasets reflecting general-domain language tasks.
@@
 Specifically, a Bradley-Terry reward model was trained on human-generated preference data obtained through pairwise comparisons of model outputs.
-The reward model evaluates the policy model’s responses based on helpfulness, accuracy, coherence, and alignment with human intent and outputs an outcome reward value for each Bai et al.
-(2022).
+The reward model evaluates the policy model’s responses based on helpfulness, accuracy, coherence, and alignment with human intent and outputs an outcome reward value for each  (Bai et al., 2022).
 Reward Model
@@
 This diversity promotes robustness and generalization in the responses generated by the policy model.
-Score Length Per Token Score
+Score Length Per Token
 REINFORCE++ 46.7 832 0.0561
-REINFORCE++ Baseline 44.2 834 0.0530
+REINFORCE++-Baseline 44.2 834 0.0530
 GRPO 46.8 860 0.0544
@@
 Experimental Results
-We use Chat-Arena-Hard Li et al.
-(2024) to evaluate our models.
+We use Chat-Arena-Hard (Li et al., 2024) to evaluate our models.
 As illustrated in Table 1, GRPO achieves a slightly superior overall score of 46.8 compared to 46.7 for REINFORCE++.
@@
 Consequently, when evaluating performance on a per-token basis, REINFORCE++ outperforms GRPO, achieving a higher per-token score of 0.0561 compared to GRPO’s 0.0544.
-The table suggests that REINFORCE++ provides more efficient outputs despite the preference bias of the evaluation model on the length.
+The table suggests that REINFORCE++ provides more efficient outputs despite the preference bias of the evaluation model on the length (Dubois et al., 2024).
 Figure 2: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and KL Divergence with Outcome Reward Model.
@@
 The evaluation results are presented in Table 2.
-GSM8K MATH HumanEval MBPP Avg.
+GSM8K MATH HumanRval MBPP Avg.
+Base Model (before training) 95.83 68.80 82.71 82.2 82.39
 REINFORCE++ 96.21 75.20 85.98 84.39 85.45
@@
 Figure 6: Comparison between GRPO and REINFORCE++ on smoothed Training Reward and Response Length with Rule Reward Model. The $x$ axis represents the training steps.
+4.2.4 Agent RL Setting: Tool-Integrated Reasoning and Agent Experiments
+Experimental Setup
+Our experiments are conducted within the framework established by Mai et al.
+(2025), which features a zero-shot agent environment designed for large language models to tackle mathematical problems.
+This environment allows the models to invoke external tools, such as Python execution, without the need for supervised demonstrations of tool usage.
+To ensure a fair comparison, we adhere to the training and evaluation protocols of the ZeroTIR environment.
+The backbone model utilized in our experiments is Qwen 2.5 Base 7B, which is initialized from pre-trained weights to maintain consistency across trials.
+The training process is facilitated by OpenRLHF Hu et al.
+(2025a), utilizing datasets from ORZ Hu et al.
+(2025b) and DAPO Yu et al.
+(2025).
+We assess the performance of the models on several established mathematical benchmarks, including AIME 2024 Jia , AIME 2025 math ai , HMMT FEB 2024/2025, and CMIMC.
+The evaluation metric employed is the average@32, consistent with the methodology outlined in Mai et al.
+(2025).
+Experimental Result
+The results of our experiments, summarized in Table 5, indicate that the REINFORCE++-baseline algorithm consistently outperforms both GRPO and PPO across all benchmarks, achieving the highest average accuracy of 24.10.
+In comparison, GRPO achieved an average accuracy of 22.58, while PPO lagged behind with an average accuracy of 21.85.
+This demonstrates the effectiveness of global batch advantage normalization in complex scenarios.
+For detailed scores across individual benchmarks, please refer to Appendix Tool-Integrated Reasoning and Agent Experiments.
+5 Limitations
+While REINFORCE++ demonstrates strong performance across diverse practical scenarios, it is important to contextualize its limitations with a balanced perspective on its strengths.
+In iid (independent and identically distributed) settings, REINFORCE++ performs comparably to GRPO—often with improved stability.
+However, it does not clearly outperform GRPO in these structured environments, where GRPO’s mature optimization remains competitive.
+This reflects REINFORCE++’s intent as a stable alternative, not a dramatic performance leap, in iid contexts.
+A more notable constraint arises from the absence of a critic model, which limits its performance ceiling relative to methods like PPO.
+PPO refines advantage estimates via a value function, while REINFORCE++ relies solely on reward-to-go.
+This simplicity aids stability but hinders precision in tasks with sparse or delayed rewards.
+Furthermore, due to compute limitations, we cannot conclusively assess REINFORCE++ at larger training scales (e.g., 3k+ steps or 8k+ long chains of thought), where its final behavior remains uncertain.
+For example, Magistral (Rastogi et al., 2025) trained with a method quiet similar to REINFORCE++-baseline achieve stronger results under extended regimes, suggesting potential headroom for improvement.
+Finally, REINFORCE++ builds on the REINFORCE family with baseline normalization.
+While it delivers strong empirical gains (especially in out-of-distribution (OOD) scenarios), it does not offer foundational theoretical innovations.
+Its contributions lie in practical refinements targeting known limitations of earlier variants.
+These observations underscore that REINFORCE++ excels as a robust, stable solution—particularly in OOD contexts—while its role in iid or critic-heavy tasks should be weighed against specific needs.
 6 Conclusion
-This paper presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences.
-Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global batch mean reward as a baseline to prevent overfitting and instability.
+In this paper, we presented REINFORCE++, a novel critic-free RLHF approach designed to efficiently align large language models (LLMs) with human preferences.
+Unlike prior methods that use prompt-specific baselines, REINFORCE++ employs a global advantage normalization to prevent overfitting and instability.
 Extensive experiments demonstrated that our algorithm achieves strong performance and improved computational efficiency across multiple RLHF scenarios, including Bradley-Terry, rule-based reward models, and long-form Chain-of-Thought (CoT) tasks, showing superior generalization capabilities.
-Future work includes exploring adaptive normalization techniques, advanced variance reduction methods, and extending REINFORCE++ beyond RLHF settings.
-Appendix A Algorithm Details
-REINFORCE++-baseline.
-For ease of use and reproducibility, we implement our proposed method within the OpenRLHF framework (https://github.com/OpenRLHF/OpenRLHF).
-Our experiments above showed that training models using multiple generated responses per prompt can further improve performance.
-Accordingly, we introduce a variant called REINFORCE++-Baseline, which integrates REINFORCE++ with multiple-response generation.
-Specifically, we sample multiple responses for each prompt and compute their average reward as the baseline to reshape the rewards.
-The advantage for each prompt-response pair is then defined as:
-$Adv_{q,o_{t}}=R_{q,o_{t}}-\text{mean}_{group}(R_{q,o_{t}})$
-$Adv_{q,o_{t}}^{norm}=\frac{Adv_{q,o_{t}}-\text{mean}_{batch}(Adv_{q,o_{t}})}{\text{std}_{batch}(Adv_{q,o_{t}})}$
-where $group$ denotes the generated responses corresponding to the same prompt.
-This baseline calculation is similar to the approach taken by GRPO; however, we deliberately move the standard deviation ($std$) of the group normalization to the global batch normalization.
-In practice, we observed that including the standard deviation introduces training instability and exacerbates overfitting to easier prompts in the RLHF process.
-For the KL constraint, REINFORCE++-baseline uses $k_{2}$ instead of the $k_{3}$ divergence approximation method in GRPO.
-For more details, please refer to the next paragraph.
-KL Penalty Design.
-In the design of REINFORCE-based RLHF methods, the choice of KL divergence estimation significantly influences algorithmic stability and accuracy.
-The widely utilized $k_{3}$ estimation in GRPO, defined by the loss function:
-$\mathcal{L}_{k_{3}}=\mathbb{E}_{x\sim D,y\sim\pi_{\theta_{\text{old}}}(\cdot|x)}\left(\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-\log\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta}(a_{t}|s_{t})}-1\right)$ (12)
-Which involve inherent biases due to its approximation nature. Specifically, the gradient of $k_{3}$ estimation approximates the gradient of the $k_{2}$ estimation through linearization:
+Future work includes exploring adaptive normalization techniques, and extending REINFORCE++ beyond RLHF settings.
+Appendix A Proof: The GRPO Advantage Estimator is Biased
+A.1 Assumptions and Settings
+We observe $N$ rewards $r_{i}$ for a prompt, and assume the true baseline is $\theta$, such that
+$r_{i}=\theta+\epsilon_{i},\quad\epsilon_{i}\sim\mathcal{N}(0,\sigma^{2}),\quad i=1,\dots,N,$
+with all advantage value $\epsilon_{i}$ independent. Define
+$\bar{\epsilon}=\frac{1}{N}\sum_{j=1}^{N}\epsilon_{j},\quad D=\sqrt{\frac{1}{N}\sum_{j=1}^{N}(\epsilon_{j}-\bar{\epsilon})^{2}},\quad A_{i}=\frac{\epsilon_{i}-\bar{\epsilon}}{D}.$
+We will prove the following:
+Theorem 1.
+For any finite $N\geq 2$, the advantage estimator $A_{i}$ is biased:
+$\mathbb{E}[A_{i}\mid\epsilon_{i}]\neq\epsilon_{i}.$
+Proof.
+Step 1: Numerator Bias
+We rewrite the numerator:
+$\epsilon_{i}-\bar{\epsilon}=\left(1-\frac{1}{N}\right)\epsilon_{i}-\frac{1}{N}\sum_{j\neq i}\epsilon_{j}.$
+Since the $\epsilon_{j}$ for $j\neq i$ are zero-mean and independent of $\epsilon_{i}$,
+$\mathbb{E}[\epsilon_{i}-\bar{\epsilon}\mid\epsilon_{i}]=\left(1-\frac{1}{N}\right)\epsilon_{i}.$
+Step 2: Denominator Depends on $\epsilon_{i}$
+(a) Compute $\mathbb{E}[D^{2}\mid\epsilon_{i}]$:
+By definition,
+$D^{2}=\frac{1}{N}\sum_{j=1}^{N}(\epsilon_{j}-\bar{\epsilon})^{2}=\frac{1}{N}\sum_{j=1}^{N}\epsilon_{j}^{2}-\bar{\epsilon}^{2}.$
+Since
+$\bar{\epsilon}=\frac{1}{N}\left(\epsilon_{i}+\sum_{j\neq i}\epsilon_{j}\right),$
+and conditioning on $\epsilon_{i}$ keeps the $\epsilon_{j}$, $j\neq i$, i.i.d. $\mathcal{N}(0,\sigma^{2})$, we obtain:
+$\mathbb{E}\left[\sum_{j=1}^{N}\epsilon_{j}^{2}\mid\epsilon_{i}\right]=\epsilon_{i}^{2}+(N-1)\sigma^{2},$
+$\mathbb{E}[\bar{\epsilon}^{2}\mid\epsilon_{i}]$ $=\frac{1}{N^{2}}\mathbb{E}\left[\left(\epsilon_{i}+\sum_{j\neq i}\epsilon_{j}\right)^{2}\Big{|}\epsilon_{i}\right]$
+$=\frac{1}{N^{2}}\left(\epsilon_{i}^{2}+2\epsilon_{i}\cdot\underbrace{\mathbb{E}\left[\sum_{j\neq i}\epsilon_{j}\right]}_{0}+\mathbb{E}\left[\left(\sum_{j\neq i}\epsilon_{j}\right)^{2}\right]\right)$
+$=\frac{\epsilon_{i}^{2}+(N-1)\sigma^{2}}{N^{2}}.$
+Subtracting:
+$\mathbb{E}[D^{2}\mid\epsilon_{i}]$ $=\frac{1}{N}(\epsilon_{i}^{2}+(N-1)\sigma^{2})-\frac{\epsilon_{i}^{2}+(N-1)\sigma^{2}}{N^{2}}$
+$=\underbrace{\frac{(N-1)^{2}}{N^{2}}\sigma^{2}}_{\alpha}+\underbrace{\frac{N-1}{N^{2}}}_{\beta}\epsilon_{i}^{2}=\alpha+\beta\epsilon_{i}^{2}.$ (15)
+(b) $g(\epsilon_{i})$ is Not Constant:
+Let $g(\epsilon_{i})=\mathbb{E}\left[1/D\mid\epsilon_{i}\right]$ and $\mu(\epsilon_{i})=\mathbb{E}[D^{2}\mid\epsilon_{i}]=\alpha+\beta\epsilon_{i}^{2}$.
+Using Taylor expansion of $f(x)=1/\sqrt{x}$ around $x_{0}=\mu(\epsilon_{i})$:
+$f(x)=\frac{1}{\sqrt{x_{0}}}-\frac{1}{2}\frac{x-x_{0}}{x_{0}^{3/2}}+\frac{3}{8}\frac{(x-x_{0})^{2}}{x_{0}^{5/2}}+O((x-x_{0})^{3})$
+Taking conditional expectation:
+$g(\epsilon_{i})$ $=\mathbb{E}[1/D\mid\epsilon_{i}]=\mathbb{E}\left[\frac{1}{\sqrt{D^{2}}}\mid\epsilon_{i}\right]=\mathbb{E}\left[f(D^{2})\mid\epsilon_{i}\right],\quad\text{where }f(x)=\frac{1}{\sqrt{x}}$
+$\approx f(\mu(\epsilon_{i}))+f^{\prime}(\mu(\epsilon_{i}))\cdot\mathbb{E}[D^{2}-\mu(\epsilon_{i})\mid\epsilon_{i}]+\frac{f^{\prime\prime}(\mu(\epsilon_{i}))}{2}\cdot\mathbb{E}[(D^{2}-\mu(\epsilon_{i}))^{2}\mid\epsilon_{i}]$
+$=\frac{1}{\sqrt{\mu(\epsilon_{i})}}-\frac{1}{2\mu(\epsilon_{i})^{3/2}}\cdot\underbrace{\mathbb{E}[D^{2}-\mu(\epsilon_{i})\mid\epsilon_{i}]}_{=\,0}+\frac{3}{8\mu(\epsilon_{i})^{5/2}}\cdot\text{Var}(D^{2}\mid\epsilon_{i})$
+$=\frac{1}{\sqrt{\mu(\epsilon_{i})}}+\frac{3}{8}\cdot\frac{\text{Var}(D^{2}\mid\epsilon_{i})}{\mu(\epsilon_{i})^{5/2}}$
+Since $\mu(\epsilon_{i})=\alpha+\beta\epsilon_{i}^{2}$ with $\beta>0$, the first term alone shows that $g(\epsilon_{i})$ depends on $\epsilon_{i}^{2}$ and hence is not constant.
+Step 3: Putting It Together
+Decomposing $A_{i}$,
+$A_{i}=\frac{\epsilon_{i}-\bar{\epsilon}}{D}=\left(1-\frac{1}{N}\right)\frac{\epsilon_{i}}{D}-\left(\frac{1}{N}\sum_{j\neq i}\epsilon_{j}\right)\cdot\frac{1}{D}.$
+[…]
+$\mathbb{E}\left[\left(\frac{-1}{N}\sum_{j\neq i}\epsilon_{j}\right)\cdot\frac{1}{D}\Big{|}\epsilon_{i}\right]=0.$
+It follows that
+$\mathbb{E}[A_{i}\mid\epsilon_{i}]=\left(1-\frac{1}{N}\right)\epsilon_{i}\cdot\mathbb{E}\left[\frac{1}{D}\mid\epsilon_{i}\right]=\left(1-\frac{1}{N}\right)\epsilon_{i}\cdot g(\epsilon_{i}).$
+Step 4: Concluding the Bias
+If $A_{i}$ were unbiased, we would have:
+$\left(1-\frac{1}{N}\right)g(\epsilon_{i})\equiv 1\quad\Rightarrow\quad g(\epsilon_{i})\equiv\frac{N}{N-1},$
+which contradicts Step 2. Therefore, for any finite $N\geq 2$,
+$\mathbb{E}[A_{i}\mid\epsilon_{i}]\neq\epsilon_{i}.$
+Hence $A_{i}$ is a biased estimator. ∎
+A.2 Why Use Global Batch Normalization?
+We observe that as $N\to\infty$, the denominator of the estimator converges to the constant $\sigma$, and the bias in the numerator vanishes.
+This insight underlies our decision to adopt global batch normalization in both REINFORCE++ and REINFORCE++-Baseline, as the global batch size is typically much larger than the group batch size in practice.
+Appendix B Algorithm Details
+B.1 KL Penalty Design
+In the design of REINFORCE-based RLHF methods, the choice of KL-divergence estimation significantly influences algorithmic stability and accuracy.
+We briefly introducing the widely utilized KL estimators,
+$\mathcal{L}_{k_{2}}$ $=\mathbb{E}_{s\sim D,\,a\sim\pi_{\theta_{\text{old}}}(\cdot|s)}\left(\frac{1}{2}(-\log x)^{2}\right)$
+$\mathcal{L}_{k_{3}}$ $=\mathbb{E}_{s\sim D,\,a\sim\pi_{\theta_{\text{old}}}(\cdot|s)}\bigl{(}(x-1)-\log x\bigr{)}\qquad$
+$\text{where}\ x$ $=\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$
+The widely utilized $k_{3}$ estimator in GRPO involves inherent biases due to its approximation nature. First, we directly provide the gradients of k2 and k3 with respect to the policy:
 $k_{2}$ Gradient $:\log x\cdot\nabla_{\theta}\log\pi_{\theta},$ (16)
 $k_{3}$ Gradient $:(x-1)\cdot\nabla_{\theta}\log\pi_{\theta},$ (17)
-where $x=\frac{\pi_{\text{ref}}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$.
-Although the $k_{3}$ gradient aligns closely with the $k_{2}$ gradient when the policy is close to the reference ($x\approx 1$), substantial deviations result in two key deficiencies:
-  1. 1.
-    Bias: As the current policy diverges significantly from the reference, particularly when $\pi_{\text{ref}}\gg\pi_{\theta_{\text{old}}}$, the linear approximation error expands nonlinearly, causing considerable biases.
-  2. 2.
-    Asymmetry: The function $(x-1)$ exhibits asymmetric behavior depending on whether $\pi_{\theta_{\text{old}}}$ surpasses or falls below $\pi_{\text{ref}}$, further complicating convergence properties.
+It is not hard to see that under the vanilla policy gradient ($r\cdot\nabla_{\theta}\log\pi_{\theta}$), the gradient with respect to k2 is equivalent to the gradient when placing the logarithmic KL term in the reward $r$, therefore, k2 is unbiased.
+When performing a Taylor expansion in the policy neighborhood ($x\approx 1$), $\log x\approx x-1$, At this time, the k3 gradient forms a linear approximation of the k2 gradient.
+However, this approximation has two key flaws:
+  • Bias: When the policy significantly deviates from the reference policy ($x$ is far from 1, especially in the later stages of training when $\pi_{\text{ref}}$ is far from $\pi_{\theta_{\text{old}}}$, particularly when $\pi_{\text{ref}}>>\pi_{\theta_{\text{old}}}$), the approximation error grows nonlinearly.
+  • Asymmetry: The response characteristics of $x-1$ are asymmetric for $\pi_{\theta_{\text{old}}}>\pi_{\text{ref}}$ and $\pi_{\theta_{\text{old}}}<\pi_{\text{ref}}$.Therefore, "k3 estimation as a loss function" is merely an approximation of "k2 estimation as a loss function".
 Therefore, despite its computational simplicity, $k_{3}$ estimation introduces biases and increased variance, suggesting it is not strictly superior to the theoretically unbiased $k_{2}$ estimation.
 Experimental evidence consistently reveals that employing $k_{3}$ estimation in GRPO results in higher variance fluctuations than $k_{2}$ estimation.
-Further discussion can be found in Liu (2025).
-Relationship with PPO.
-REINFORCE++ shares certain similarities with PPO in its formulation.
-Specifically, when PPO adopts Generalized Advantage Estimation (GAE) parameters with $\lambda=1$ and discount factor $\gamma=1$, REINFORCE++ reduces to PPO without the critic network and employs global batch normalization as the baseline.
-Mathematically, this relationship is expressed as follows:
-$\text{GAE}(\lambda=1,\gamma=1)=\sum_{l=0}^{\infty}r_{t+l}-V(s_{t}),$ (15)
-where $V(s_{t})$ denotes the estimated value function at state $s_{t}$.
-By removing the critic network, the REINFORCE++ algorithm effectively eliminates the $V(s_{t})$ term.
-At the same time, the introduction of global batch normalization further stabilizes training by reducing variance and preventing overfitting to individual prompts.
-Other Implementation Details
+Further discussion can be found in (Liu, 2025).
+B.2 Implementation Tricks
 Mini-Batch Updates
@@
   • Scaling: Applies appropriate scaling factors for numerical stability during updates.
-Appendix B Acknowledgements
-  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the early draft of the paper.
+Appendix C Tool-Integrated Reasoning and Agent Experiments
+In this section, we present a comprehensive evaluation of the effectiveness of various reinforcement learning algorithms in the context of tool-augmented reasoning.
+Our experiments are conducted within the framework established by Mai et al.
+(2025), which features a zero-shot agent environment designed for large language models to tackle mathematical problems.
+This environment allows the models to invoke external tools, such as Python execution, without the need for supervised demonstrations of tool usage.
+To ensure a fair comparison, we adhere to the training and evaluation protocols of the ZeroTIR environment.
+The backbone model utilized in our experiments is Qwen 2.5 Base 7B, which is initialized from pre-trained weights to maintain consistency across trials.
+The training process is facilitated by OpenRLHF Hu et al.
+(2025a), utilizing datasets from ORZ Hu et al.
+(2025b) and DAPO Yu et al.
+(2025).
+We assess the performance of the models on several established mathematical benchmarks, including AIME 2024 Jia , AIME 2025 math ai , HMMT FEB 2024/2025, and CMIMC.
+The evaluation metric employed is the average@32, consistent with the methodology outlined in Mai et al.
+(2025).
+[…]
+The results of our experiments, summarized in Table 5, indicate that the REINFORCE++-baseline algorithm consistently outperforms both GRPO and PPO across all benchmarks, achieving the highest average accuracy.
+This finding aligns with previous research by Mai et al.
+(2025), which highlighted the stability and effectiveness of the REINFORCE++-baseline and global advantage normalization in tool-augmented environments.
+Algorithm AIME 24 AIME 25 HMMT 2025 HMMT 2024 CMIMC Avg
+GRPO 31.66 21.87 16.97 17.70 24.68 22.58
+PPO 30.20 21.66 15.00 18.43 23.95 21.85
+RF++-baseline 30.83 27.18 17.91 18.95 25.62 24.10
+Table 5: Performance Comparison Across Benchmarks (average@32).
+Appendix D Best Practices
+This paper introduces two algorithms: REINFORCE++ and REINFORCE++-Baseline, with a key focus on selecting the appropriate algorithm based on task conditions.
+Empirical evidence from the open-source community suggests that REINFORCE++-Baseline is particularly effective for sample filtering or more complex scenarios.
+For instance, in the multi-turn tool-calling setting discussed earlier, a high proportion of void (i.e., non-informative) samples can destabilize training.
+In such cases, incorporating a baseline—by subtracting the intra-group mean reward—significantly improves training stability by effectively filtering out void samples.
+Moreover, automatic reward reshaping reduces the need for designing complex reward structures.
+For example, REINFORCE++-Baseline supports both 0/1 and -1/1 reward schemes, while REINFORCE++ performs best with symmetric rewards, such as -1/1 in RLVR tasks.
+In contrast, for tasks where obtaining multiple reward signals for distinct responses conditioned on the same intermediate state is challenging—such as training with Process-Supervised Reward Models (PRMs) or online real-time sampling—we recommend using REINFORCE++.
+Appendix E Acknowledgements
+  • Jian Hu: Developed the core ideas and implemented the algorithms for REINFORCE++ and REINFORCE++-baseline, as well as contributed to the proof of GRPO advantage estimator.
   • Jason Klein Liu: Implemented the experimental code, fine-tuned hyperparameters, wrote the paper, and provided GPU resources.
-  • Wei Shen: Led the paper writing and supervised the main experiments.
+  • Haotian Xu: Conducted comparative experiments between reinforce++-baseline and complex tool calling as well as agent scenarios.
+  • Wei Shen: Led the paper writing and designed and supervised the main experiments.